Circular No. 5844

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1992bt IN NGC 3780 A. V. Filippenko, T. Matheson, and B. Leibundgut, University of Cal-

ifornia at Berkeley, report that a CCD spectrum (range 310-1000 nm) obtained on Aug. 13 UT with the Shane 3-m reflector at Lick Observatory shows that SN 1992bt (cf. IAUC 5841) is of type II. The object is in its nebular phase, with strong Hα emission and weaker lines of [O I] 630-nm, [Ca II] 730-nm, and the Ca II near-infrared triplet.

GRO J0422+32 P. Garnavich, Dominion Astrophysical Observatory, writes: “CCD

photometry obtained at the Manastash Ridge Observatory of the University of Washington gives the following preliminary V magnitudes: Aug. 10.476 UT, V = 17.8 ± 0.1; 11.470, 15.83 ± 0.04; 12.455, 15.68 ± 0.03; 13.490, 15.32 ± 0.03.

C. Chevalier and S. A. Ilovaisky, Observatoire de Haute-Provence, report: “CCD photometry of the optical counterpart of the soft x-ray transient GRO J0422+32 was obtained with the OHP 1.2-m telescope from July 25 to Aug. 16. When the object was faint, we find V = 19.0 on July 26.1 UT and V = 19.2 on July 28.1. Monitoring in the Cousins I band started in August, yielding the following preliminary magnitudes: Aug. 7.1, 18.1; 8.1, 17.6; 9.1, 17.8; 12.1, 15.0; 13.1, 14.8; 14.1, 14.5; 16.1, 14.6. Observations on Aug. 16 yield V = 15.3, B–V = +0.5, V–I_{c} = +0.7.”

SUPERNOVA 1993J IN NGC 3031 G. F. Lawrence, A. Paulson, C. Mason, C. Butenhoff, and R. D. Gehrz

report the following infrared photometry of SN 1993J obtained at the University of Minnesota 0.76-m telescope (+ InSb photometer; standard infrared photometric techniques employed with a 27′′ beamsize and typically a 44′′ north/south throw; calibration was against IRTF standard stars): Apr. 3.4 UT, J = 10.91 ± 0.05, H = 10.66 ± 0.05, K = 10.36 ± 0.08; 21.2, 10.01 ± 0.05, 10.00 ± 0.05, 9.80 ± 0.05; 22.2, 10.05 ± 0.03, 9.97 ± 0.03, 9.85 ± 0.03; May 13.3, 10.92 ± 0.02, 10.48 ± 0.03, 10.48 ± 0.03; June 22.2, 12.45 ± 0.09, 11.69 ± 0.06, 11.20 ± 0.09; July 7.2, 13.03 ± 0.15, 12.13 ± 0.07, 11.73 ± 0.12; 15.2, 12.92 ± 0.09, 12.51 ± 0.06, 11.89 ± 0.10. Note that on IAUC 5757 the authors incorrectly reported the K magnitude for Apr. 3.4 as 10.48.

1993 August 17 Daniel W. E. Green